OpenStudy (anonymous):

a. Solve the differential equation y prime equals the product of 4 times x and the square root of the quantity 1 minus y squared. b. Explain why the initial value problem y prime equals the product of 4 times x and the square root of the quantity 1 minus y squared with y(0) = 4 does not have a solution.

OpenStudy (anonymous):

https://gyazo.com/fa06c87f78388c31263201846039dc26

terenzreignz (terenzreignz):

So... go ahead, show me you aren't clueless XD

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Not clueless but somewhat stuck

OpenStudy (anonymous):

I remember this but still quite lost

OpenStudy (anonymous):

How would I solve

OpenStudy (anonymous):

\[1-y^2=e^{2x^2+c}\]

OpenStudy (anonymous):

For C?

OpenStudy (anonymous):

Need a refreshing mint

terenzreignz (terenzreignz):

What in the world... haha To be honest, I haven't even started trying to solve it myself :> Hang on...

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

There's a way for sure, just can't seem to remember

terenzreignz (terenzreignz):

\[\Large \frac{dy}{dx}=4x\sqrt{1-y^2}\]

terenzreignz (terenzreignz):

\[\Large \frac {dy}{\sqrt{1-y^2}}=4xdx\] You got this far?

OpenStudy (anonymous):

Yup and integrated.

terenzreignz (terenzreignz):

And you got...?

OpenStudy (anonymous):

\[\ln(\sqrt{1-y^2})=2x^2+C\]

terenzreignz (terenzreignz):

Ahh. No.

terenzreignz (terenzreignz):

Not even close :P

terenzreignz (terenzreignz):

Humour me and try differentiating \[\Large \ln \sqrt{1-y^2}\] You'll see you won't get to the original \[\Large \frac1{\sqrt{1-y^2}}\]

OpenStudy (anonymous):

Is that not the original?

OpenStudy (anonymous):

\[\frac{ y }{ y^2-1 }\]

terenzreignz (terenzreignz):

No. Chain rule.

terenzreignz (terenzreignz):

And the original is \[\Large \frac1{\sqrt{1-y^2}}\]

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

It is a property of inverse sin

OpenStudy (anonymous):

Am I right or right?

terenzreignz (terenzreignz):

Finally ^^ \[\Large \sin^{-1}(y) = 2x^2 + C\]

OpenStudy (anonymous):

Now what?

terenzreignz (terenzreignz):

Well, the second part of the question. Show that the initial value y(0) = 4 has no solution.

terenzreignz (terenzreignz):

Should be easy enough?

OpenStudy (anonymous):

So wait....that's the end of the first part? No way to make it y=?

OpenStudy (anonymous):

I honestly don't think so but sometimes you poop magic out

OpenStudy (anonymous):

Magic that makes sense.

terenzreignz (terenzreignz):

Well, you can certainly have \[\Large y = \sin[2x^2 + C]\]

terenzreignz (terenzreignz):

It doesn't really matter. Now, do the second part. :D

OpenStudy (anonymous):

I like that better. Second part it is

OpenStudy (anonymous):

You get

OpenStudy (anonymous):

\[4=\sin(c)\]

OpenStudy (anonymous):

Which can't be solved.

terenzreignz (terenzreignz):

Sehr gut ^^ Very good.

OpenStudy (anonymous):

wut?

terenzreignz (terenzreignz):

Sehr gut is "very good" in German :D

OpenStudy (anonymous):

TIL

terenzreignz (terenzreignz):

Are we done here? :D